OpenStudy (anonymous):

One factor of 10x^3+51x^2+3x-10 is (x+5). what are the other two factors? (my answer: (x-1/4) and (x+1/2)

OpenStudy (aum):

Incorrect. Do a synthetic division of 10x^3+51x^2+3x-10 by (x+5). Then factor the quadratic.

OpenStudy (anonymous):

I got this, buthow do you factor this p(x) (x+5) (10x^2+x-2)?

OpenStudy (aum):

10x^2 + x - 2 10 * -2 = -20 what two factors of -20 add to +1? 5 and -4 10x^2 + x - 2 10x^2 + 5x - 4x - 2 5x(2x + 1) - 2(2x + 1) (5x-2)(2x+1)

OpenStudy (aum):

10x^3+51x^2+3x-10 = (x+5)(5x-2)(2x+1)

OpenStudy (anonymous):

thank you sooo much! i got it!

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

can I ask you another question?

OpenStudy (aum):

go ahead.